OpenStudy (anonymous):

A set of data items is normally distributed with a mean of 60 and a standard deviation of 4. Convert 63 to a z-score. Round your answer to the nearest hundredth as needed.

OpenStudy (anonymous):

what topic is this?

OpenStudy (anonymous):

Normal Distributions and Z Scores

OpenStudy (anonymous):

stats..., @Dany612

jimthompson5910 (jim_thompson5910):

z = (x - mu)/sigma z = (63-60)/4 z = 3/4 z = 0.75

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome